i'm having trouble finding the answer to this problem 1/2(x-8)=x-16 what are the steps i should take?
Answer by richwmiller(17219)

multiply both sides by 2
x-8=2x-32
add 32 to both sides
x+24=2x
subtract x from both sides
24=x
check
1/2(24-8)=24-16
1/2(16)=8
8=8
ok